OpenStudy (anonymous):

Which expressions would complete this equation so that it has infinitely many solutions? 8 + 2(8x – 6) = Choose exactly two answers that are correct. A. 2(4x + 7) i think its <--- and B. 9x – 10 C. 16x – 4 this one <-- :) D. 4(4x – 1)

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

i put 8 + 2 (8x - 6) = 2(4x + 7)

OpenStudy (anonymous):

Let's rewrite the first term a little: 8+2(8x-6)=16x-4. Rewrite the other four as well: A: 2(4x+7)=8x+14 B: 9x-10 C: 16x-4 D: 4(4x-1)=16x-4 Okay so firstly we can see that C is the same as our first equation so we know this is an answer because 16x-4=16x-4 works for any x. We also worked out that D is equal to 16x-4 so this is your second answer :)

OpenStudy (anonymous):

then i ended up with 8 + 2(8x - 6) = 2(4x +7) 8 + 16x - 12 = 8x + 7 16x - 4

OpenStudy (anonymous):

O.O whaaaaaaaaaaaaaaaaaaaaa?

OpenStudy (anonymous):

lol, which bit don't you get?

OpenStudy (anonymous):

sso you say the answer is C and D?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

well can you break it down how you broke it down before

OpenStudy (anonymous):

Do you understand what I did to A, B, C and D?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'm just simplifying the terms, so for A: 2(4x + 7)=8x+14

OpenStudy (anonymous):

Are you with me so far?

OpenStudy (anonymous):

uh so your pluging in 8 + 2(8x - 6)?

OpenStudy (anonymous):

with the answer's?

OpenStudy (anonymous):

Slow down a bit :) Okay so the question specifies that we're looking for answers with infinitely many solutions, so we want equations like x=x, do you agree?

OpenStudy (anonymous):

now i wrote on the paper 8 + 2(8x-6) = 16x-4 8+ 16x -12 6x - 4 6x-4 = 6x -4?

OpenStudy (anonymous):

where looking for x = x you say?

OpenStudy (anonymous):

Equations of that form, yes. y=y, z=z it doesn't matter, all we want is when both sides are the same because this is true regardless of the value of x we pick. If we have x=x, then this is true for 1: 1=1, 2: 2=2, 3: 3=3 and so on. It's also the same for 16x-4=16x-4 :) As our initial equation simplifies down to 16x-4, we're looking to find two answers that simplify down to this as well - which, in this case, is C and D :)

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

HOW YOU KNOW YOU LOOKIN FOR A EQUAL?

OpenStudy (anonymous):

FOR IT TO BE EQUAL

OpenStudy (anonymous):

We want infinite solutions, there are ONLY infinite solutions when the equation is of the form x=x.

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

SOO

OpenStudy (anonymous):

HOW IS IT C AND D WHAT DID YOU SOLVE JUST BREAK DOWN THE NUMBERS ON HOW YOU GOT IT TO BE C AND B

OpenStudy (anonymous):

I MEAN C AND D

OpenStudy (anonymous):

TOM!?

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

-.- THANX FOR UR HELP II GUESS

OpenStudy (anonymous):

I'm still here, but I've got work of my own to do as well so I might come and go for a few minutes.

OpenStudy (anonymous):

Do you agree our initial expression is 16x-4?

OpenStudy (anonymous):

This is the same as C, so C is an answer.

OpenStudy (anonymous):

how you use 16x - 4 and you pick C and D? o.o

OpenStudy (anonymous):

Simplify 8 + 2(8x – 6)

OpenStudy (anonymous):

i get 16x - 4

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

Now look for A-D which is 16x-4.

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

now i get it

OpenStudy (anonymous):

Thank god, I've run out of ways to explain it haha. Glad you understand.

OpenStudy (anonymous):

lol thats all you had to say THANKS!

OpenStudy (anonymous):

sorry for not understanding i would have quite if i was you tbh

OpenStudy (anonymous):

kk bye

OpenStudy (anonymous):

No need to be sorry, I was in your position 10 years ago so it's nice to pass on the favour :)

OpenStudy (anonymous):

lol k :)